# Insensitive edge solitons in non-Hermitian topological lattices

### v1

@@ Abstract
Finally, we present results beyond the weakly nonlinear regime and its perturbative analysis. To do so we perform numerical continuation like in Fig. 2 for all NMs with $\gamma=0.6$ and $s=0.6$, in region (II), toward $I\rightarrow\infty$. The results are shown in Fig. 5. It displays persistent NESs with $E\approx 0$ (red lines) even at large intensity. In fact, we see that as we increases the nonlinearity from $I=0$, the famlies of NESs originating from the TM ($k=0$) at constant $E\approx 0$, and the one emerging from the SM with $k=-1$ (grey line) of growing $E$, inexorably comes closes around $I=2.5$ without crossing. We find at this point the characters of the two families are exchanged [left inset of Fig. 5]. Consequently, the family of NESs emerging from the $k=-1$ SM has constants $E\approx 0$ for growing $I$ pass $I=2.5$. Further, we project on the NM basis the high-amplitude NES at $I=2.75$ and find weak couplings of the TM with its surrounding, like what is seen for weak amplitude NESs, right-inset of Fig. 5. Furthermore, we find a cascade of such transitions between the consecutive families of nonlinear modes emerging from the SMs with $k=-1$ and $-2$, $k=-2$ and $-3$, etc.
@@
In summary, we have investigated the spectral sensitivity of edge solitons arising from the topological modes within a nonreciprocal Su-Schrieffer-Heeger chain featuring local Kerr nonlinearity. We found in large enough lattices, such a family of nonreciprocal non-edge solitons (NESs) with zero-energy without requiring chiral symmetry, as usually the case [54], for e.g., when implementing complex nonlinear couplings [67, 16]. Furthermore, we anticipate that our findings can be extended to other classes of nonlinear localized modes, such as breathers [68, 69]. Consequently, our work may open new ways to manipulate waves in topological systems for applications in sensing, lasing and information processing.
@@ Acknowledgements.
V.A and B.M.M acknowledge the support of the program Etoiles Montantes en Pays de la Loire and the EU H2020 ERC StG “NASA” Grant Agreement No. 101077954.

### v2

@@ Abstract
To go one step further, we present results beyond the weakly nonlinear regime and its perturbative analysis. To do so we perform numerical continuation like in Fig. 2 for all eigenmodes with $\gamma=0.6$ and $s=0.6$, in region (II), toward $I\rightarrow\infty$. The results are shown in Fig. 5. It displays persistent NES with $E\approx 0$ (red lines) even at large intensity. In fact, we see that as we increases the nonlinearity from $I=0$, the famlies of NES originating from the TM ($l=0$) at constant $E\approx 0$, and the one emerging from the bulk with $l=-1$ (grey line) of growing $E$, inexorably comes closes around $I=2.5$ without crossing. We find at this point the characters of the two families are exchanged [left inset of Fig. 5]. Consequently, the family of NES emerging from the $l=-1$ bulk mode has constants $E\approx 0$ for growing $I$ pass $I=2.5$. Further, we project on the eigenmode basis the high-amplitude NES at $I=2.75$ and find weak couplings of the TM with its surrounding, like what is seen for weak amplitude NES, right-inset of Fig. 5. Furthermore, a cascade of such transitions appears between consecutive families of nonlinear modes emerging from the bulk modes (e.g., $l=-1$ kicks $-2$, $l=-2$ kicks $-3$, etc). It would be worthwhile to explore the mechanisms at the origin of these zero-energy highly nonlinear NES, using more precise methods of bifurcation like normal form theory [73].
@@
In summary, we have investigated the spectral sensitivity of edge solitons arising from the topological modes within a nonreciprocal Su-Schrieffer-Heeger chain featuring local Kerr nonlinearity. We found a family of nonreciprocal edge solitons (NES) with zero-energy without requiring chiral or sublattice symmetries, as usually the case [57], for e.g., when implementing complex nonlinear couplings [74, 16]. Furthermore, we anticipate that our findings can be extended to other classes of nonlinear localized modes, such as breathers [75, 76]. Consequently, our work may open new ways to manipulate waves in topological systems for applications in sensing, lasing and information processing.
@@ Acknowledgements.
V.A and B.M.M acknowledge the support of the program Etoiles Montantes en Pays de la Loire and the EU H2020 ERC StG “NASA” Grant Agreement No. 101077954. The authors would like to express their gratitude to the anonymous reviewers for their comments and suggestions, which helped improve the clarity of this manuscript.